which equation describes the line that contains the points (-1, -8) and (-9, -21)?
1. (y+8)=13/8(x+1)
2. y=3x-8
3. 2y-18x=5

Answer by stanbon(75887)

which equation describes the line that contains the points
(-1, -8) and (-9, -21)?
---
slope = (-21--8)/(-9--1) = -13/-8 = 13/8
----
Ans: Only y+8 = (13/8)x+1 has that slope.
Cheers,
Stan H.
==============
1. (y+8)=13/8(x+1)
2. y=3x-8
3. 2y-18x=5
-------